OpenStudy (anonymous):

The sum of the measure of the interior angles of a polygon is 1080°. What polygon is it?

OpenStudy (eliasperez191):

i think it is a...... awe crap i dont remeber let me check ok

hartnn (hartnn):

for a polygon with side 'n' the sum of measures of interior angles is \(\large (n-2)\times 180\) so just find n from \(1080 = (n-2)\times 180\)

OpenStudy (anonymous):

i don't think i understand what your telling me. how does this determine what kind it is?

hartnn (hartnn):

based on number of sides, we can tell what kind of polygon... like if we get n=7, we say its septagon (or heptagon ) if we get n =8, we say its octagon and so on

OpenStudy (princeharryyy):

sum of interior angles of a polygon is given by (n-2)*180 where n is the number of sides of the polygon , here the total sum of interior angles is 1080 degree so the equation becomes (n-2)*180 = 1080 which gives n=8 and hence the answer is octagon which has 8 sides.